OpenStudy (anonymous):

2x2 + 13x + 15 2x2 + 27x + 13

OpenStudy (anonymous):

Factor?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\Delta=13^{2}-4*2*15, x _{1,2}=(-b \pm \Delta)/2\] result for the first ecuation

OpenStudy (anonymous):

-b=-13

OpenStudy (anonymous):

First 1: a=2, b=13, c=15----->a(c)= 2(15)= 30=3(10) 2x^2+10x+3x+15 2x(x+5)+3(x+5) (2x+3)(x+5) Same idea for #2 2x^2+x+26x+13 X(2x+1)+13(2x+1) (X+13)(2x+1)

OpenStudy (anonymous):

\[x _{1,2}=(-13\pm \Delta)/2^{2}\]

OpenStudy (anonymous):

thank you both :)

OpenStudy (anonymous):

You r wlcm!